# Height filtrations and base loci on flag bundles over a curve

### v1

@@ Abstract.
Let $k$ be an algebraically closed field of characteristic zero. Let $C/k$ be a projective smooth curve with function field $K=k(C)$ and $G/k$ be a connected reductive group. Let $F$ be a principal $G$-bundle on $C$. Let $P\subseteq G$ be a parabolic subgroup and $\lambda:P\longrightarrow G$ be a strictly anti-dominant character. Then $F/P\longrightarrow C$ is a flag bundle and $\mathcal{L}_{\lambda}=F\times_{P}k_{\lambda}$ on $F/P$ is a relatively ample line bundle.
Let $h_{\mathcal{L}_{\lambda}}$ be the height function on $X=(F/P)_{K}$ induced by $\mathcal{L}_{\lambda}$. We compute its height filtration and successive minima: the height filtration is given by a Bruhat decomposition and successive minima are given by slopes.
An interesting application is that, in this case, Zhang’s inequality on successive minima can be enhanced into an equality. This is proved from the convex geometry point of view.
Let $f\in N^{1}(F/P)$ be the numerical class of a vertical fiber. We also compute the augmented base loci $\mathrm{B}_{+}(\mathcal{L}_{\lambda}-tf)$ for any $t\in\mathbb{R}$ and it turns out that they are almost the same as the height filtration, which accords with the philosophy in Arakelov geometry that the positivity of the line bundle $\mathcal{L}_{\lambda}$ is equivalent to the positivity of its induced height function $h_{\mathcal{L}_{\lambda}}$. As a corollary, we compute the $k$-th movable cones of flag bundles over curves for all $k$.
@@ 1. Introduction
The main purpose of this paper is to give an explicit description of the height function and its related topics on a flag bundle over a curve. In this paper, all fields are of characteristic zero and all curves are projective smooth. By variety we mean an integral seperated scheme of finite type over a field.
@@ 1.1. Height filtration and successive minima
We start by giving a reminder on height functions in Arakelov theory and the relevant history. Let $K$ be either a number field or $K=k(C)$ where $C$ is a projective smooth curve over a field $k$. Let $X/K$ be a projective variety of dimension $d$ and $\overline{L}$ be an adelic line bundle on $X$. These data induce an Arakelov height function $h_{\overline{L}}$ on $X$ ([27], see also [25, §9] for a survey). A typical case is the geometric height, which is the one we concern in this article. Here we give a definition.
If $K=k(C)$, let $\mathcal{X}\longrightarrow C$ be a projective flat morphism with generic fiber $X\longrightarrow\operatorname{Spec}(K)$ and $\mathcal{L}$ be a line bundle on $\mathcal{X}$ with $\mathcal{L}_{K}\cong L$. The data $(\mathcal{X},\mathcal{L})$ define an adelic line bundle $\overline{L}$ and the height function $h_{\overline{L}}$ is given by
@@
We also denote this by $h_{\mathcal{L}}$ if there is no ambiguity. If $K$ is a number field, the height function can be defined similarly by arithmetic intersection theory.
Assume $L$ is ample. Let $Z_{t}(X,h_{\overline{L}})=\text{Zariski closure of}\;\big{\{}x\in X(\overline{K}):h_{\overline{L}}(x)<t\big{\}}$. Note that
  • $t\longmapsto Z_{t}(X,h_{\overline{L}})$ is an increasing filtration of Zariski closed subsets.
  • $Z_{t}(X,h_{\overline{L}})=X$ when $t\gg 0$ and $Z_{t}(X,h_{\overline{L}})=\emptyset$ when $t\ll 0$.
Since Zariski topology is Noetherian, the filtration $Z_{t}(X,h_{\overline{L}})$ gives a finite filtration $X_{0}=X\supsetneq X_{1}\supsetneq X_{2}\supsetneq\cdots\supsetneq X_{r}=\emptyset$, the height filtration. Its jumping points $\zeta_{i}(X,h_{\overline{L}})=\inf\big{\{}t:Z_{t}(X,h_{\overline{L}})=X_{i-1}\big{\}}$ are called the successive minima.
@@
Note that our definition of successive minima are slightly different with Zhang [27]: Zhang considers only dimension jumps $e_{i}(X,h_{\overline{L}})=\inf\big{\{}t:\dim Z_{t}(X,h_{\overline{L}})\geq d-i+1\big{\}}$, $i=1,\dots,d+1$. In the following, $e_{i}(X,h_{\overline{L}})$ will be called Zhang’s successive minima to avoid ambiguity.
The first minimum $\zeta_{1}(X,h_{\overline{L}})=e_{1}(X,h_{\overline{L}})=\inf\big{\{}t:Z_{t}(X,h_{\overline{L}})=X\big{\}}$ is of particular interest, it is called the essential minimum and also denoted by $\zeta_{\text{ess}}(X,h_{\overline{L}})$. This invariant is the protagonist of the story of the Bogomolov conjecture, which in number field case says that a closed subvariety $X$ of an abelian variety $A$ has essential minimum zero with respect to the Néron-Tate height on $A$, if and only if it is the translation of an abelian subvariety by a torsion point, we refer the reader to [26, 24] for more details. The height filtration induced by the Néron-Tate height is $A\supsetneq\emptyset$ with the unique jumping point $\zeta=0$. This allows the equidistribution of small points to happen, and consequently implies the Bogomolov conjecture.
@@
The height filtration (of toric height functions) on toric varieties is completely determined in [10]. An important example is on the projective space $\mathbb{P}^{n}_{\mathbb{Q}}$. Consider the height function induced by the universal bundle $\mathcal{O}_{\mathbb{P}^{n}_{\mathbb{Z}}}(1)$ equipped with the Fubini-Study metric $\|\cdot\|_{\text{FS}}$. The height filtration is given by successively deleting torus orbits, that is, $\mathbb{P}^{n}_{\mathbb{Q}}\supsetneq\bigcup_{i}\big{\{}x_{i}=0\big{\}}\supsetneq\bigcup_{i,j}\big{\{}x_{i}=0,x_{j}=0\big{\}}\supsetneq\cdots\supsetneq\bigcup_{i}\big{\{}x_{i}\neq 0\text{ and $x_{j}=0$ for all $j\neq i$}\big{\}}\supsetneq\emptyset$ with successive minima $\frac{1}{2}\log(k+1)$, $k=0,1,\dots,n$.
In this article, we consider another class of varieties with group actions: flag varieties. Let $k$ be an algebraic closed field, $G/k$ be a connected reductive group, $C/k$ be a curve, and $F$ be a principal $G$-bundle on $C$ with canonical reduction $F_{Q}$ to $Q$. This defines a linear functional ${\mathrm{deg}}(F_{Q})$ on $X(T)_{\mathbb{Q}}$. Let $P\subseteq G$ be a parabolic subgroup and $\lambda:P\longrightarrow G$ be a strictly anti-dominant character. Then $F/P\longrightarrow C$ is a flag bundle and $\mathcal{L}_{\lambda}=F\times_{P}k_{\lambda}$ on $F/P$ is a relatively ample line bundle. Let $X=(F/P)_{K}$ and $h_{\mathcal{L}_{\lambda}}$ be the height function on $X$ induced by $(F/P,\mathcal{L}_{\lambda})$. We compute its height filtration and successive minima.
@@
The height filtration of $h_{\mathcal{L}_{\lambda}}$ on $X$ is given by successively deleting Schubert cells $C_{w}=(F_{Q}\times_{Q}QwP/P)_{K}$ for $w\in W_{Q}\backslash W/W_{P}$, i.e.
@@
In particular, successive minima are $\zeta_{w}=\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle$ and Zhang’s successive minima are $e_{i}=\min\big{\{}\zeta_{w}:\ell(w)=\dim G/P-i+1\big{\}}$ where $\ell(w)=\max\limits_{\sigma\in W_{Q}wW_{P}}\min\limits_{\tau\in\sigma W_{P}}\ell(\tau)$.
1.2. Zhang’s inequality
Let $X/K$ be a projective variety of dimension $d$, and $\overline{L}$ be a relative semipositive adelic line bundle with the underlying line bundle $L$ being ample. The height of $X$ is defined to be $h_{\overline{L}}(X)=\frac{\overline{L}^{d+1}}{(d+1)L^{n}}$. In [27], Zhang gives the following inequality:
$e_{1}\geq h_{\overline{L}}(X)\geq\frac{1}{d+1}\sum_{i=1}^{d+1}e_{i}.$
In the case of toric varieties, Zhang’s successive minima can be intepreted by convex geometry as shown in [9, 10]. Moreover, Zhang’s inequality can be proved directly [10, Theorem C] as an integral inequality of a convex function on a polytope. This can be partially generalized to arbitrary projective varieties by the theory of Okounkov bodies [19] and concave transforms [8, 2].
In this article, we carry on this spirit in the setting of flag bundles $F/P$. Given a strictly anti-dominant character $\lambda:P\longrightarrow\mathbb{G}_{m}$, representation theorists could associate to $\lambda$ a string polytope $\Delta_{\lambda}$, a weight polytope $\mathcal{P}_{\lambda}$ and a weight map $p:\Delta_{\lambda}\longrightarrow\mathcal{P}_{\lambda}$. Let $M_{\lambda}=G\times_{P}k_{\lambda}$ be the line bundle on $G/P$ induced by $\lambda$. Then integral points in $\Delta_{\lambda}$ parametrizes a weight basis of $H^{0}(G/P,M_{\lambda})$ and $p:\Delta_{\lambda}\longrightarrow\mathcal{P}_{\lambda}$ sends an integral point to the weight of the vector parametrized by it [18, §2.2].
Since the weight polytope is the convex hull of the Weyl orbit $W\lambda$, it follows immediately from Theorem A that successive minima of $h_{\mathcal{L}_{\lambda}}$ are the value of ${\mathrm{deg}}(F_{Q})$ at vertices of $\mathcal{P}_{\lambda}$. We will also prove
$h_{\mathcal{L}_{\lambda}}(X)=\frac{1}{\operatorname{Vol}(\Delta_{\lambda})}\int_{\Delta_{\lambda}}{\mathrm{deg}}(F_{Q})\circ p\;\mathrm{d}\mu.$
An interesting application is that, Zhang’s inequality in this case can be enhanced into an equality, in the sense that
@@ 1.2. Zhang’s inequality
Let $q:W/W_{P}\longrightarrow W_{Q}\backslash W/W_{P}$ be the canoniacl projection and for $w\in W_{Q}\backslash W/W_{P}$ define $m_{w}=\#q^{-1}(w)$. Then
$h_{\mathcal{L}_{\lambda}}(X)=\frac{1}{\#W/W_{P}}\sum_{w\in W_{Q}\backslash W/W_{P}}m_{w}\zeta_{w}.$
This follows from a simple computation of integration on the string polytopes: in fact for any linear functional $f:X(T)_{\mathbb{R}}\longrightarrow\mathbb{R}$, we have
We remark that it is certainly not possible to enhance Zhang’s inequality into an equality in general. Consider the height function on $\mathbb{P}^{n}_{\mathbb{Q}}$ induced by the the Hermitian line bundle $\big{(}\mathcal{O}_{\mathbb{P}^{n}_{\mathbb{Z}}}(1),\|\cdot\|_{\text{FS}}\big{)}$. The variety height is $\frac{1}{2}\sum_{i=1}^{n}\sum_{j=1}^{i}\frac{1}{j}$ [7], which is a rational number while the successive minima are irrational numbers $\tfrac{1}{2}\log(k+1)$, $k=0,1,\dots,n$.
1.3. Base loci
It is a general philosophy in Arakelov geometry that the positivity of $\overline{L}$ are equivalent to the positivity of its induced height function $h_{\overline{L}}$. The study of essential minimum and pseudo-effectivity provides an evidence [10, 2, 22]. Note that if the underlying line bundle $L$ is big, then $\overline{L}$ is big $\Longleftrightarrow$ $\zeta_{\mathrm{ess}}(h_{\overline{L}})>0$.
On the other hand, the theory of augmented base loci $\mathrm{B}_{+}(\cdot)$ is well-developed [13, 14] in the study of positivity in algebraic geometry. Now back to the case $\mathcal{X}=F/P$, let $f$ be a line bundle given by a vertical fiber. Recall that it is a classical result that $\mathrm{B}_{+}(\mathcal{L}_{\lambda}-tf)$ is the union of closed subvariety over which the restricted volume is 00 (see §4.1). In particular, if $\mathcal{Y}\subseteq\mathcal{X}$ is a closed subvariety such that $(\mathcal{L}_{\lambda}-tf)|_{\mathcal{Y}}$ is not big on $\mathcal{Y}$, then $\mathcal{Y}\subseteq\mathrm{B}_{+}(\mathcal{L}_{\lambda}-tf)$.
For $w\in W_{Q}\backslash W/W_{P}$ such that $\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle\leq t$, our calculation of $h_{\mathcal{L}_{\lambda}}$ gives that the essential minimum $\zeta_{\mathrm{ess}}(h_{(\mathcal{L}_{\lambda}-tf)|_{\mathcal{X}_{\omega}}})$ on the Schubert variety $\mathcal{X}_{w}=F_{Q}\times_{Q}\overline{QwP}/P$ is no bigger than 00, hence $(\mathcal{L}_{\lambda}-tf)|_{\mathcal{X}_{\omega}}$ is not big by the arithmetic bigness criterion mentioned above. We thus obtain that
$\bigcup_{\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle\leq t}\mathcal{X}_{w}\subseteq\bigcup_{\mathcal{Y}\subseteq\mathcal{X},(\mathcal{L}_{\lambda}-tf)|_{\mathcal{Y}}\text{ not big}}\mathcal{Y}\subseteq\mathrm{B}_{+}(\mathcal{L}_{\lambda}-tf)$
where $\mathcal{Y}$ runs over all closed subvariety of $\mathcal{X}$. Moreover, we prove that both inclusions are actually equality:
@@ 1.3. Base loci
The augmented base locus of $\mathcal{L}_{\lambda}-tf$ is given by
$\mathrm{B}_{+}(\mathcal{L}_{\lambda}-tf)=\mathcal{X}\Bigg{\backslash}\coprod_{\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle>t}\mathcal{C}_{w}=\coprod_{\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle\leq t}\mathcal{C}_{w}$
where $\mathcal{C}_{w}=F_{Q}\times_{Q}QwP/P$.
In particular, comparing this with Theorem A, we see that $\mathrm{B}_{+}(\mathcal{L}_{\lambda}-tf)_{K}=Z_{t}$ for all $t\in\mathbb{R}$, except for the critical points $t=\zeta_{w}$ at which $B_{+}(\mathcal{L}_{\lambda}-\zeta_{w}f)$ is upper continuous while $Z_{\zeta_{w}}$ is lower continuous.
1.4. Movable cones
Recall that for a projective variety $Y$ of dimension $d$, the k-th movable cones $\operatorname{Mov}^{k}(Y)$, $k=1,2,\dots,d$ are defined as
$\operatorname{Mov}^{k}(Y)=\big{\{}\alpha\in N^{1}(Y)_{\mathbb{R}}:\mathrm{B}_{+}(\alpha)\text{ has codimension $\geq k$}\big{\}}.$
$\operatorname{Mov}^{k}(Y)$ are open cones in $N^{1}(Y)_{\mathbb{R}}$ and we have obvious inclusions
$\operatorname{Mov}^{d}\subseteq\operatorname{Mov}^{d-1}\subseteq\cdots\subseteq\operatorname{Mov}^{1}.$
Note that $\operatorname{Mov}^{1}$ is the big cone and $\operatorname{Mov}^{d}$ is the ample cone.
Let $k$ be an algebraically closed field, $C/k$ be a curve and $E$ be a vector bundle on $C$. Note that $N^{1}(\mathbb{P}(E))_{\mathbb{R}}$ is two dimensional, generated by $f$ and $\xi=\mathcal{O}_{\mathbb{P}(E)}(1)$. The ample cone is given by extremal ray $f$ and $\xi-\mu_{\min(E)}f$, and the big cone is given by extremal ray $f$ and $\xi-\mu_{\max(E)}f$. The former is a theorem of Miyaoka [21] and the latter can be found in [15]. Fulger and Lehmann computed $\operatorname{Mov}^{2}(\mathbb{P}(E))$ in [16]. In [4], the big cone of the Grassmann bundle of $E$ is computed and in [6], the ample cone of flag bundles of $E$ is computed.
As a direct consequence of Theorem C, we compute the $k$-th movable cones of flag bundles $F/P$ over a curve $C$ for all $k$. Since any element in $N^{1}(F/P)_{\mathbb{R}}$ can be written as $\mathcal{L}_{\lambda}-tf$, we may define two functions $\langle\alpha^{\vee},\cdot\rangle$ and $\langle{\mathrm{deg}}(F_{Q}),w\cdot\rangle$ on $N^{1}(F/P)_{\mathbb{R}}$ as
These functions are well-defined, and

### v2

# Arakelov geometry on flag varieties over function fields and related topics

@@ Abstract.
Let $k$ be an algebraically closed field of characteristic zero. Let $G$ be a connected reductive group over $k$, $P\subseteq G$ be a parabolic subgroup and $\lambda:P\longrightarrow G$ be a strictly anti-dominant character. Let $C$ be a projective smooth curve over $k$ with function field $K=k(C)$ and $F$ be a principal $G$-bundle on $C$. Then $F/P\longrightarrow C$ is a flag bundle and $\mathcal{L}_{\lambda}=F\times_{P}k_{\lambda}$ on $F/P$ is a relatively ample line bundle.
We compute the height filtration, successive minima, and the Boucksom-Chen concave transform of the height function $h_{\mathcal{L}_{\lambda}}:X(\overline{K})\longrightarrow\mathbb{R}$ over the flag variety $X=(F/P)_{K}$. An interesting application is that the height of $X$ equals to a weighted average of successive minima, and one may view this as a refinement of Zhang’s inequality of successive minima.
Let $f\in N^{1}(F/P)$ be the numerical class of a vertical fiber. We compute the augmented base loci $\mathrm{B}_{+}(\mathcal{L}_{\lambda}-tf)$ for any $t\in\mathbb{R}$, and it turns out that they are almost the same as the height filtration. As a corollary, we compute the $k$-th movable cones of flag bundles over curves for all $k$.
@@ 1.1. Height filtration and successive minima
We start by giving a reminder on height functions in Arakelov theory. Let $K$ be either a number field or $K=k(C)$ where $C$ is a projective smooth curve over a field $k$. Let $X$ be a projective variety of dimension $d$ over $K$ and $\overline{L}$ be an adelic line bundle on $X$. These data induce an Arakelov height function $h_{\overline{L}}$ on $X$ (see [24, §9] for a survey). A typical case is the geometric height, which is the one we concern in this article. Here we give the definition.
If $K=k(C)$, consider a projective flat morphism $\mathcal{X}\longrightarrow C$ with the generic fiber $X\longrightarrow\operatorname{Spec}(K)$ and a line bundle $\mathcal{L}$ on $\mathcal{X}$ with $\mathcal{L}_{K}\simeq L$. The data $(\mathcal{X},\mathcal{L})$ define an adelic line bundle $\overline{L}$ and the height function $h_{\overline{L}}$ is given by
@@
We also denote this by $h_{\mathcal{L}}$ if there is no ambiguity. If $K$ is a number field, the height function can be defined similarly via arithmetic intersection theory.
@@
For any $t\in\mathbb{R}$, let $Z_{t}\subseteq X$ be the Zariski closure of the set $\big{\{}x\in X(\overline{K}):h_{\overline{L}}(x)<t\big{\}}$. Let’s call $\big{\{}Z_{t}:t\in\mathbb{R}\big{\}}$ the height filtration, and call its jumping points the successive minima. Note that our definition of successive minima are slightly different with Zhang [25]. Zhang considers only dimension jumps $e_{i}=\inf\big{\{}t:\dim Z_{t}\geq d-i+1\big{\}}$, $i=1,\dots,d+1$. In this article, $e_{i}$ will be called the successive minima of Zhang to avoid ambiguity.
Determining the height filtration is generally difficult and only two cases are known in the literature:
  • The Néron-Tate height function on abelian varieties $A$ over $K$. Let $L$ be an ample symmetric line bundle over $A$ endowed with the unique adelic metric such that $[2]^{*}\overline{L}=2^{\dim A}\overline{L}$. The height filtration of the induced Néron-Tate height function $\widehat{h}_{L}$ is $A\supsetneq\emptyset$ and the jumping point is $\zeta=0$;
@@
  • The toric height function on toric varieties [8]. They are given by successively deleting torus orbits. An important example is the projective space $\mathbb{P}^{n}_{\mathbb{Q}}$ and the height function induced by the universal bundle $\mathcal{O}_{\mathbb{P}^{n}_{\mathbb{Z}}}(1)$ equipped with the Fubini-Study metric $\|\cdot\|_{\text{FS}}$. The height filtration on $\mathbb{P}^{n}_{\mathbb{Q}}$ is given by
    $\quad\quad\quad\mathbb{P}^{n}_{\mathbb{Q}}\supsetneq\bigcup_{i}\big{\{}x_{i}=0\big{\}}\supsetneq\bigcup_{i,j}\big{\{}x_{i}=0,x_{j}=0\big{\}}\supsetneq\cdots\supsetneq\emptyset$
    with successive minima $\frac{1}{2}\log(k+1)$, $k=0,1,\dots,n$.
Our first result (Theorem A below) provides a new case where height filtrations can be explicitly computed. Let $k$ be an algebraically closed field of characteristic zero and $C$ be a projective smooth curve over $k$ with function field $K=k(C)$. Let $G$ be a connected reductive group over $k$, $P\subseteq G$ be a parabolic subgroup and $\lambda:P\longrightarrow\mathbb{G}_{m}$ be a strictly antidominant character. Let $F$ be a principal $G$-bundle over $C$ and $\mathcal{X}=F/P$ with generic fiber $X=(F/P)_{K}$. Let $\mathcal{L}_{\lambda}=F\times_{P}k_{\lambda}$ and $h_{\mathcal{L}_{\lambda}}:X(\overline{K})\longrightarrow\mathbb{R}$ be the induced height function.
${X=(F/P)_{K}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{X}=F/P\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{Spek}(K)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${C}$
Let $F_{Q}$ be the canonical reduction of $F$ to $Q$, and ${\mathrm{deg}}(F_{Q})\in X(T)^{\vee}_{\mathbb{Q}}$ be the induced cocharacter. Let $W,W_{P}$ and $W_{Q}$ be the Weyl group of $G,L_{P}$ and $L_{Q}$. For $w\in W_{Q}\backslash W/W_{P}$, let $C_{w}=(F_{Q}\times_{Q}QwP/P)_{K}\subseteq X$ be the corresponding Schubert cell. The numbers $\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle$ are well-defined.
@@
For any $t\in\mathbb{R}$, let $Z_{t}\subseteq X$ be the Zariski closure of the set $\big{\{}x\in X(\overline{K}):h_{\mathcal{L}_{\lambda}}(x)<t\big{\}}$. Then
@@
In particular, the filtration $\big{\{}Z_{t}:t\in\mathbb{R}\big{\}}$ is finite (while in general height filtrations could be infinite) and is given by successively deleting Schubert cells $C_{w}$ with jumping points $\zeta_{w}=\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle$, $w\in W_{Q}\backslash W/W_{P}$.
1.2. String polytopes and the Boucksom-Chen concave transform
Attached to a projective variety $X$ over $K$ and an adelic line bundle $\overline{L}$ on $X$ with $L$ ample and $\overline{L}$ semipositive, we have the Okounkov body $\Delta$ and the Boucksom-Chen concave transform $\varphi:\Delta\longrightarrow\mathbb{R}$ thanks to [17, 18, 6, 1]. The maximum of $\varphi$ equals to the essential minimum of $h_{\overline{L}}$, the minimum of $\varphi$ equals to the absolute minimum of $h_{\overline{L}}$ and the integral mean of $\varphi$ equals to the height of $X$.
In toric cases, they can be expressed as summation of local roof functions [7]. But in general, the Boucksom-Chen concave transform is difficult to compute. Our second result (Theorem B) provides a new case where the Boucksom-Chen concave transform can be explicitly computed.
To state the result, keep the notations in Theorem A. Let $w_{0}\in W$ be the longest element and fix a reduced decomposition $\underline{w_{0}}=s_{i_{1}}s_{i_{2}}\cdots s_{i_{N}}$, $N=\dim G/B$. Then we have the string valuation $\nu$ and the string polytope $\Delta_{\lambda}\subseteq\mathbb{R}^{N}$ [19, 3], which is a polytope of dimension $d=\dim G/P$. Let $\mathcal{P}_{\lambda}$ be the weight polytope and $p:\Delta_{\lambda}\longrightarrow\mathcal{P}_{\lambda}$ be the weight map. The Okounkov body of $(X,L)$ (with respect to a specific valuation) is the string polytope $\Delta_{\lambda}$.
@@ 1.2. String polytopes and the Boucksom-Chen concave transform
The Boucksom-Chen concave transform of ${\mathcal{L}_{\lambda}}$ is $\Delta_{\lambda}\longrightarrow\mathbb{R},x\longmapsto\langle{\mathrm{deg}}(F_{Q}),p(x)\rangle$.
Let $\mu$ be the Lebesgue measure on the real span of $\Delta_{\lambda}$, normalized with respect to $\mathbb{Z}^{N}$. As we mentioned above, the variety height is given by the integral mean of the Boucksom-Chen concave transform [6]. Thus a direct corollary is
$h_{\mathcal{L}_{\lambda}}(X)=\frac{\mathcal{L}_{\lambda}^{\dim\mathcal{X}}}{\dim\mathcal{X}\cdot L^{\dim X}}=\frac{1}{\operatorname{Vol}(\Delta_{\lambda})}\int_{\Delta_{\lambda}}\langle{\mathrm{deg}}(F_{Q}),p(x)\rangle\,\mathrm{d}\mu(x).$
The integration can be computed via Weyl symmetry.
@@
Let $q:W/W_{P}\longrightarrow W_{Q}\backslash W/W_{P}$ be the canoniacl projection. For $w\in W_{Q}\backslash W/W_{P}$, let $m_{w}=\#q^{-1}(w)$. Then
$h_{\mathcal{L}_{\lambda}}(X)=\frac{1}{\#W/W_{P}}\sum_{w\in W_{Q}\backslash W/W_{P}}m_{w}\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle.$
Note that $\zeta_{w}=\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle$ in Theorem C are successive minima of the height function $h_{\mathcal{L}_{\lambda}}$ on $X$. Thus Theorem C builds an equality between the variety height and a weighted average of the successve mimina. One may recognize this as a refinement of Zhang’s inequality of successive minima [25]
$e_{1}\geq h_{\mathcal{L}_{\lambda}}(X)\geq\frac{1}{d+1}\sum_{i=1}^{d+1}e_{i}.$
into an equality. Here $e_{i}$ is the the successive minima of Zhang.
We remark that it is not possible to build such an equality in general. Consider the height function on $\mathbb{P}^{n}_{\mathbb{Q}}$ induced by the Hermitian line bundle $\big{(}\mathcal{O}_{\mathbb{P}^{n}_{\mathbb{Z}}}(1),\|\cdot\|_{\text{FS}}\big{)}$. The variety height is a rational number $\frac{1}{2}\sum_{i=1}^{n}\sum_{j=1}^{i}\frac{1}{j}$ while the successive minima are irrational numbers $\tfrac{1}{2}\log(k+1)$, $k=0,1,\dots,n$.
1.3. Base loci and movable cones
It is a general philosophy in Arakelov geometry that the positivity of $\overline{L}$ is equivalent to the positivity of the induced height function $h_{\overline{L}}$. The study of essential minimum and pseudo-effectivity [8, 1, 22] provides an evidence.
Our third result (Theorem D below) provides a new evidence to support the above mentioned philosophy. Keep the notations in Theorem A. Let $f\in N^{1}(F/P)$ be the class of a vertical fiber and $\mathrm{B}_{+}(\mathcal{L}_{\lambda}-tf)$ be the augmented base locus of $\mathcal{L}_{\lambda}-tf$ for $t\in\mathbb{R}$. They measure the positivity of the line bundle $\mathcal{L}_{\lambda}$.